OpenStudy (anonymous):

Write the vertex form of the equation of the parabola which has a vertex of (0, 0) and contains the point (2, –2). help please.

OpenStudy (abdullahm):

Welcome to OpenStudy! @museee Do you know what the vertex form is?

OpenStudy (anonymous):

thank you . and its a(x-h)^2+k, right ?

OpenStudy (abdullahm):

correct, where (h,k) is the vertex

OpenStudy (abdullahm):

So first of all, lets plug in the vertex into this equation. :)

OpenStudy (anonymous):

but how do i find that ? im sorry im really terrible at math

OpenStudy (abdullahm):

vertex is (0,0) in the form of (h,k) that means h=0 and k=0

OpenStudy (abdullahm):

so if we plug that into the vertex form we get: \(\sf\Large y=a(x-h)^2+k\) \(\sf\Large y=a(x-0)^2+0\) \(\sf\Large y=a(x)^2\)

OpenStudy (abdullahm):

Now plug in the points (2,-2) to solve for 'a'